OpenStudy (anonymous):

The corner section of seats in a stadium contains 2 seats in the front row, 4 seats in the second row, 6 seats in the third row, etc. Each row has 2 more seats than the row in front of it. There are 2n seats in the nth row. How many total seats are in a corner section with n rows of seats? A)n2 + 2 B)n2 + n C)n2 + 2n D)n2 + 2n + 1

OpenStudy (anonymous):

n^2 + n

OpenStudy (anonymous):

we have (2 + 4 + 6 + ... + 2n) = 2(1+2+3+...+n) sum to n of natural numbers = (1/2)(n(n+1)) multiply by 2 to get B